OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4).

OpenStudy (mertsj):

Can you find the slope of this line?

OpenStudy (anonymous):

\[y=4/3x+8\]

OpenStudy (anonymous):

Kelly is correct :)

OpenStudy (anonymous):

parallel slope = 4/3 4 = (4/3) (-3) + b ->b = 4 + 4 = 8 Thus y = 4x/3 + 8